# Learning to Drive Anywhere with Model-Based Reannotation

### v1

@@ 5.4 Evaluating MBRA on effectively using crowd-sourced data
In the final experiment, we aim to assess the generalization capabilities of MBRA model. To this end, we deploy the short-horizon navigation policy on robots in diverse environments across 6 countries: USA, Mexico, China, Mauritius, Costa Rica, and Brazil, as shown in the appendix. In total, we collect 24 topological graphs and evaluate each trajectory. To the best of our knowledge, we are the first to conduct a global evaluation for visual navigation. We evaluate multi-step VPT and MBRA model trained with and without the FrodoBots-2k dataset. Findings are summarized in Table 6, with more fine-grained results in the appendix. MBRA model as short-horizon goal image-conditioned navigation policy had better performance for both goal reaching and subgoal coverage than multi-step VPT.
Figure 6: Comparing multi-step VPT and MBRA.
Comparing multi-step VPT and MBRA The key differences between multi-step VPT and MBRA are highlighted in Fig. 6. If the noise in the dataset comes from a Gaussian distribution, as shown in Fig. 6[a], the action labels in the data are expected to be inconsistent for observations along the GT trajectories. And it is expected that the GT trajectories itself are noisy due to data collection by non-expert teleoperators. Imitating such a noisy trajectories as shown in Fig. 6[b] is impractical because it will be heavily skewed by intermediate inconsistent noisy actions, leading to incorrect reannotations. In contrast, MBRA is more robust to noisy data because it prioritizes the final goal pose, which is typically further from the individual positions and therefore can leverage all FrodoBots-2k dataset to train MBRA, leading better reannotation for entire FrodoBots-2k dataset.
6 Conclusion
Limitations: Our MBRA approach to reannotating noisy crowd-sourced data and action-free in-the-wild videos in the long-horizon navigation setting works well but leaves room for improvement. In the model-based approach, we may sometimes generate unreasonable actions because of inaccuracies in the robot model. While we find the model-based approach to generally outperform the imitation-based relabeler (multi-step VPT), it does require some strong conditions on the model itself that could prove difficult to translate to more complex tasks like manipulation. One axis of future improvement is developing a more accurate differentiable model by incorporating more accurate 3D geometry, environment semantics, and dynamic object behaviors, such as pedestrian behavior [37]. It would also be helpful to consider not only goal reaching but also to incorporate humans’ preferences into the objective design, particularly when navigating in crowds or in settings where semantic conventions are important (e.g., not driving on grass when it is inappropriate). While our model inherits some semantic behaviors (like staying on paths) from the tendencies exhibited by the human operators in the data, such preferences are not enforced explicitly.
Acknowledgments
This research was supported by Berkeley AI Research at the University of California, Berkeley and Toyota Motor North America. And, this work was partially supported by DARPA TIAMAT, ARL DCIST CRA W911NF-17-2-0181, NSF IIS-2246811, and NSF IIS-2150826. We thank Frodobots AI for providing robot hardware and computational resources for our evaluations.
Appendix A Design of extended Kalam filter
To filter the noisy GPS and compass data present in the FrodoBots-2K dataset, we apply an modified Extended Kalman Filter [40] that estimates the following state variables:
  • The robot’s current position $(x,y)\in\mathbb{R}^{2}$ in the UTM frame
  • The robot’s heading $\theta$, expressed as a pair $(\cos\theta,\sin\theta)$
  • The current compass heading bias $\gamma$
  • The current GPS bias $\delta$
We model the linear and angular velocity from instantaneous robot actions as inputs to the system, and treat measured GPS position $(x_{gps},y_{gps})$ and compass heading $\theta_{mag}$ as observed outputs with the following measurement equations:
$\begin{array}[]{cc}x_{gps}=x+\delta_{x}&y_{gps}=y+\delta_{y}\\ \cos(\theta_{mag})=\cos(\theta+\gamma)&\sin(\theta_{mag})=\sin(\theta+\gamma)\end{array}$
We model the system update as:
$x_{t+1}$ $=x_{t}+v_{t}\cos\theta_{t}\cdot\Delta t+\eta_{x}\sqrt{\Delta t}$
$y_{t+1}$ $=y_{t}+v_{t}\sin\theta_{t}\cdot\Delta t+\eta_{y}\sqrt{\Delta t}$
$\begin{pmatrix}\cos\theta_{t+1}\\ \sin\theta_{t+1}\\ \end{pmatrix}$ $=\begin{pmatrix}\cos\nu_{t}&-\sin\nu_{t}\\ \sin\nu_{t}&\cos\nu_{t}\end{pmatrix}\begin{pmatrix}\cos\theta_{t}\\ \sin\theta_{t}\end{pmatrix}$
$\quad\quad\textrm{for}\quad\nu_{t}=\omega_{t}\cdot\Delta t+\eta_{\theta}\sqrt{\Delta t}$
$\gamma_{t+1}$ $=\gamma_{t}+\eta_{\gamma}\sqrt{\Delta t}$
$\delta_{t+1}$ $=\delta_{t}+\eta_{\delta}\sqrt{\Delta t}$
The noise terms $\eta$ are modeled as i.i.d. normal. We use a forward-backward Kalman smoother that integrates estimates of two Kalman filters, one running forwards in time and another running backwards [48], which allows our estimates to make use of all available information at all timesteps.
Appendix B Model-based learning considering time delay
In training MBRA model, we introduce a novel objective design considering the variable system delay to consistently learn the robotic behavior during the delay period. In the system with $L$ steps delay, the robot has to act by the previous action commands $\{a^{s}_{i}\}_{i=-L\ldots 1}$ for $L$ steps. In other words, after $L$ steps running $\{a^{s}_{i}\}_{i=-L\ldots 1}$, the robot can act according to the action commands $a_{0}$. Hence, without taking into account the robotic behavior of $\{a^{s}_{i}\}_{i=-L\ldots 1}$ in training, the generated action command $a_{0}$ causes overshooting and/or oscillation against the target trajectories. We train MBRA by considering the past action commands $\{a^{s}_{i}\}_{i=-L\ldots 1}$ to generate more consistent action commands.
We feed the previous action commands $\{a_{i}\}_{i=-L_{max}\ldots-1}$ and a randomly selected virtual delay step $L~{}(\leq L_{max})$ into our network in addition to the visual observations. $L_{max}(=6)$ is the maximum step number of the assumed time delay. Then we calculate the following objective considering the robotic state $\{s^{s}_{i}\}_{i=-L\ldots N-1}$ for $L+N$ steps
$\min\,J_{mbl}:=\sum_{i=-L}^{N-1}(s^{ref}-s^{s}_{i})^{2},$ (5)
where we estimate the states by using the previous action commands $\{a^{s}_{i}\}_{i=-L\ldots-1}$ and the generated action commands $\{a^{s}_{i}\}_{i=0\ldots N-1}$ from our policy as $\{s^{s}_{i}\}_{i=-L\ldots N-1}=f(O_{c},\{a^{s}_{i}\}_{i=-L\ldots N-1})$. In our implementation, $s^{s}_{i}$ is defined by three components, $[\hat{p}_{i},\hat{c}_{i},\Delta a^{s}_{i}]$, to encourage the policy to smoothly move toward the target goal pose $p^{g}$ without collision. Here $\hat{p}_{i}$ is the $i$-th virtual robot pose, $\hat{c}_{i}$ is the estimated collision state at $i$-th virtual robot pose $\hat{p}_{i}$ (where zero indicates no collision), and $\Delta a^{s}_{i}$ indicates the action difference, $a^{s}_{i+1}-a^{s}_{i}$. Accordingly, we define $s^{ref}$ as [$p_{g}$, 0.0, 0.0]. By minimizing $J_{mbl}$, we can train $\pi^{s}$ while considering the behavior during the uncontrollable time delay by $\{a^{s}_{i}\}_{i=0\ldots N-1}$.
In inference, we decide $L$ depending on the system architecture and feed the generated $a_{0}$ to control the robot. In our implementation, we set $L=2$ for the ERZ and $L=0$ for other robots. A more detailed description of our model-based learning configuration is available in the original paper of ExAug [19].
Appendix C Network architecture
Figure 7 shows the network architecture of both our image-goal conditioned and goal-pose conditioned models. For short-horizon navigation policy $\pi^{s}$, we concatenate the current observation $O_{c}$ and the goal observation $O_{g}$ and generate a goal-conditioned embedding with EfficientNet-B0. In addition, we concatenate the image observation history $\{O_{i}\}_{i=-M\ldots 0}$ and generate a history embedding with EfficientNet-B0. We pass in these visual features, the system delay $L$ and the previous action commands $\{a^{s}_{i}\}_{i=-L\ldots-1}$ to a set of Transformer and fully connected MLP layers to produce a sequence of 2D poses $\{a_{i}\}_{i=0\dots N-1}$.
For the long-horizon navigation policy $\pi^{l}$, we replace the visual encoder for $O_{c}$ and $O_{g}$ with MLP layers for the local 2D goal pose $p_{g}$ on the current robot coordinate in our implementation. We calculate $p_{g}$ from the current robot pose and the target goal pose on the global coordinate. In addition, we no longer include system delay length $L$ and previous actions $\{a^{s}_{i}\}_{i=-L\ldots-1}$. Instead of considering the delay during training, we use the $L^{\textrm{th}}$ step of the output during inference, similar to [49] and [9].
Figure 7: Network architecture. In addition to the visual observations, We feed the delay step and the previous actions to consider the system delay in the model-based learning objective. For the long-horizon navigation policy, we replace the visual encoder for the current and the goal observation with the MLP layers for the goal pose.
Appendix D Action and observation space
The action space of our policy and relabeler is defined as the pose—comprising the $(x,y)$ position and yaw angle—on a 2D plane. Following the original implementation of [19], the MBRA model internally generates a sequence of linear and angular velocities, which are then integrated to produce action commands in the position space for training LogoNav.
During inference with LogoNav, we employ a PD controller to generate linear and angular velocity commands that track the generated pose commands, in accordance with the original implementations of [8, 9, 27]. Since our robotic hardware platforms including ERZ, Vizbot, and Go1 are controlled via linear and angular velocity commands (which are subsequently translated into low-level control actions such as wheel angular velocities or leg joint angles), we directly apply these commands to guide the robot toward the target goal pose. Furthermore, during inference with the MBRA model used as a goal image-conditioned navigation policy (see Sec.5.4), we directly use the linear and angular velocity commands internally generated by the model, without employing a PD controller. As visually illustrated in Figs.2 and 7, our observations consist of raw camera images. Prior to being input to the network, these images are processed using standard normalization.
Appendix E Evaluation platform
We use three robot platforms, FrodoBot “Earth Rover Zero” (ERZ), VizBot, and the quadruped robot Go1 for our evaluations.
FrodoBot “Earth Rover Zero” (ERZ): The FrodoBot “Earth Rover Zero” (ERZ), shown in Fig. 8, is a low-cast RC car used both for crowdsourced dataset collection and our navigation policy deployment. The ERZ includes a host of sensors such as front and back side cameras, GPS, an IMU unit including gyroscope, accelerometer and compass sensors, and wheel velocity sensors in all four wheels.
However, the inexpensive hardware setup of the robots and crowd-sourcing approach result in significant noise. The main factors of the noisy action labels are 1) robot inconsistencies and corresponding user adjustments, 2) low-cost GPS and IMU, 3) inevitable wheel slips during turning, 4) robot vibration during turning, and 5) system time delay. The details of the robot hardware and the reason of the noisy action labeling are shown in the appendix.
The operators can introduce action label noise by intuitively making adjustments for windy conditions, controller drift, and imbalanced wheels, which are not reflected by the visual observations of the robot. On top of that, the robot turns by using friction to slip in place, making the turning behavior highly dependent on the road condition and robot pose estimation from wheel speed sensors inaccurate. This slippage also causes significant vibration of the robot chassis that inhibits the ability to use IMU measurements for action estimation. In addition, all signals are sent between the robot and the workstation via the internet, resulting in about a 0.7-second time delay between the observations and the velocity commands. This delay must be considered when designing our policy architecture to ensure we account for its effects.
All measurements from the sensors can be accessed through the platform’s API. Linear and angular velocity commands can also be sent to the robot from user teleoperation for data collection (gaming) or from our trained policies for navigation. The rover can turn in place and lasts for five hours on a fully charged battery.
Other robotic platforms: We conduct additional evaluations with different robot hardware and systems to analyze the cross-embodiment performance of our policy. We show the overview of the VizBot [47] and the Unitree Go1 quadruped robot in Fig. 9. Different from the ERZ, we deploy our trained policy on an Nvidia Orin AGX mounted on the robot and evaluate navigation performance. Instead of using GPS, we mount a tracking camera on top of our robot for localization indoors. In addition, we use a different camera, a PCB-mounted fisheye camera, and use its image observations for inference.
Figure 8: Overview of Earth Rover Zero (ERZ) and its system. The ERZ can be controlled over a 4G internet connection for gaming and data collection and for deploying our navigation policy.
Figure 9: Overview of robotic platforms, VizBot and Go1. These robots mounts different camera from ERZ and can be controlled with an onboard robot controller on an Nvidia Orin AGX with ROS.
Appendix F Baseline methods
In our evaluation of long-horizon navigation, we use the following two baselines, NoMaD and behavior clolning (BC). For BC, we evaluate various annotation methods as the ground truth action labels to compare with our MBRA relabeler.
NoMaD [27]: We deploy the original NoMaD policy [27] for exploration and generate 30 possible trajectories. Out of these options, we select the best trajectory by measuring the distance between the last predicted position and the goal pose and selecting the minimum one to control the robot.
Behavior Cloning [9]: We train a long-horizon navigation policy on reannotated action labels by following several baseline methods instead of using our MBRA. All learning setups except annotation are same as our method.
Raw action label: As the simplest action commands, we annotate the robot trajectory with the recorded GPS and the compass readings at 1.0 Hz. To match the image frame rate, we linearly interpolate between adjacent timesteps. For training, we sample the raw robot poses for 8 steps at 3.0 Hz and transform them into the local robot coordinate frame to be used as ground truth.
Filtered action label: We give the above mentioned Extended Kalman filter (EKF) for entire FrodoBots-2k dataset to estimate the less-noisy robot pose. Similar to the raw action label, we sample the filtered pose for 8 steps at 3.0 Hz in training and transform them into the local robot coordinate.
[…]
VPT [22]: We train the inverse dynamics models to estimate the relative pose between two consecutive observations such as $p^{i}_{i+1}=f_{vpt}(O_{i},O_{i+1})$ by imitating the ground-truth relative pose in the dataset. In training, we sample 9 image flames from the current frame to the 8-step future frame at 3.0 Hz as $\{O_{i}\}_{i=0\ldots 8}$ and estimate the relative poses $\{p^{i}_{i+1}\}_{i=0\ldots 7}$ between each frame by the trained inverse dynamics model. Then we integrate the estimated relative poses $\{p^{i}_{i+1}\}_{i=0\ldots 7}$ to have the trajectories in the current local coordinate and use the local trajectories as the ground truth.
Multi-step VPT [22, 9, 28]: Since we want to annotate the actions for 8 steps, we train the model to estimate the robotic action for 8 steps such as select $O_{g}$ as the 8 step future frame from $O_{c}$ in training. Following [9], we train the robotic foundation model as multi-step VPT to estimate the robotic action to move between two frames, $O_{c}$ and $O_{g}$. Since we want to annotate the actions for 8 steps, we select $O_{g}$ as the 8 step future frame from $O_{c}$ in training. The other training setups are same as the original paper [9]. In training, we sample $O_{c}$ and $O_{g}$ (8 step future frame for $O_{c}$) and estimate the robotic action to move toward the location of $O_{g}$ and supervise its estimated action commands.
[…]
Appendix G Evaluation breakdown for long-horizon navigation policy
To evaluate the long-horizon navigation policy, we conduct 3 trials for 7 different trajectories and show the average performance in Table 2. To clarify the performance of each method, we show the average goal success rate and coverage rate for each location in Table 7. The goal success rate measures if the policy was able to reach the final goal pose along the topological graph, and the coverage rate measures how far along the trajectory the robot got. LogoNav trained with MBRA succeeds at least once in 5 environments, outperforming the other policies. This suggests that our method can leverage the diverse dataset and enables us to train the most generalized policy, which can navigate the robot toward the goal poses in more than 80 % of our evaluation environments.
Table 7: Evaluation breakdown of the long-horizon navigation using GPS localization. GS indicates the goal success rate, COV indicates the coverage rate and FRB indicates the FrodoBots-2k dataset.
NoMaD BC: filtered actions BC: visual SLAM BC: VPT BC: multi-step VPT LogoNav: MBRA
(GNM only) (GNM + FRB) (GNM + FRB) (GNM + FRB) (GNM + FRB) (GNM + FRB)
Environment GS COV GS COV GS COV GS COV GS COV GS COV
Env A 0.333 0.533 0.000 0.467 0.333 0.400 0.000 0.200 1.000 1.000 0.667 0.767
Env B 0.000 0.100 0.000 0.467 0.000 0.333 0.000 0.200 0.667 0.667 1.000 1.000
Env C 0.667 0.733 0.333 0.400 0.667 0.767 0.000 0.200 0.000 0.333 1.000 1.000
Env D 0.667 0.733 1.000 1.000 1.000 1.000 0.000 0.267 0.667 0.967 1.000 1.000
Env E 0.333 0.333 0.333 0.933 0.000 0.000 0.000 0.000 0.667 0.700 0.667 0.967
Env F 0.000 0.333 0.333 0.667 0.000 0.500 0.667 0.833 1.000 1.000 1.000 1.000
Env G 0.333 0.533 0.000 0.433 0.000 0.400 0.000 0.500 0.333 0.633 0.667 0.733
Average 0.333 0.471 0.286 0.624 0.286 0.486 0.095 0.314 0.619 0.757 0.857 0.924
Appendix H Data ablation study
[…]
Table 8: Ablation study for the FrodoBots-2k dataset. “GS” and “SC” indicate the goal success rate and the subgoal coverage rate, respectively.
Dataset multi-step VPT MBRA
@@ Appendix H Data ablation study
[…]
To evaluate the performance of each mixture in a variety of situations, we deploy the policies on four indoor trajectories and four outdoor trajectories. The distance from the initial node to the goal node is between 10.0 m and 31.0 m. As shown in Table 8, our MPC-inspired relabeler’s performance improves with the increasing fraction of data from the FrodoBots-2k. The imitation-based relabeler (multi-step VPT), on the other hand, has a minor improvement from including $1\%$ of the FrodoBot dataset, however performance deteriorates as more data is added. This is likely because the percentage of noisy action labels increases and training collapses. Therefore, MBRA is effective at providing high-quality action labels for the crowd-sourced dataset, resulting in stronger short-horizon navigation performance.
Appendix I Visualization of evaluation over the world
We deploy the short-horizon navigation policy on robots in diverse environments across 6 countries: USA, Mexico, China, Mauritius, Costa Rica, and Brazil. Figure 10 shows one sample trajectory for each country. Our evaluation environments include a variety of weather conditions (e.g. cloudy and sunny), time of day (e.g. morning, afternoon, and evening), and diverse environmental types, such as urban, suburban, and rural settings. However, since the robot operator is not around the robot, we prioritize safe experiments and do not select the time and the area with a lot of pedestrians and cars. Note that the environment selection prioritizing safety is only for the fully remote evaluation at the global scale. We conduct local experiments in human-occupied areas with many obstacles, as shown in the supplemental video.
Figure 10: Short-horizon navigation sample trajectories. We evaluate short-horizon navigation performance in a variety of environments around the world at different times and dates.
Appendix J Visualization of reannotated trajectories
Figure 11 shows an example of action labels generated by different methods in six scenes. The raw action commands (orange) do not reflect the motion expected when moving to the pose associated with the goal image. The filtered action (magenta) is more reasonable than the raw action labels; however, they are inaccurate. The trajectories from the visual SLAM(red) are also inaccurate due to the limited visual features. The VPT (cyan) is generating the action commands to move straight ahead in almost all cases. The multi-step VPT(green), which corresponds to ViNT [9], generates the action commands toward the goal pose. But the actions predicted by MBRA (blue) are most consistent with the ground truth motion that moves toward the goal image pose without collision. In case B, the blue trajectory by MBRA goes to the right side, which is slightly deviated from the goal pose to avoid collisions with the wall of the left side.
Figure 11: Reannotated action trajectories in FrodoBots-2k dataset. Our reannotated action label (green) is visually reasonable to move toward the goal location specified with the goal observation. Hand-labeled approximate true trajectories are indicated with yellow arrows.

### v3

@@ V-B Baseline Method
V-C Long-horizon Navigation Policy (LogoNav): GPS Goals
To answer Q1, we evaluate the long-horizon navigation policies trained with MBRA and several baselines. We select 7 outdoor locations and evaluate each policy 3 times for each goal. In Table II, we show the goal success rate and the coverage rate for each method. The coverage rate is the ratio of the distance reached by the robot to the distance of the target goal pose before it fails. Our policy with MBRA shows stronger performance than all baselines for both goal success rate and coverage rate. In Sec. V-E, we conduct an investigation to analyze the advantageous gap of MBRA to answer Q3.
Figure 3 shows the third-person view at the start position and the robot trajectories on a bird-eye-view map in two scenes. Our policy distilled from MBRA actions was the only one to successfully navigate to the distant goal pose in both scenes, making a sharp left turn at the start to stay on path in case A. In contrast, both NoMaD and GCP could not execute this action, failing by colliding with bushes or requiring interventions to avoid falling down stairs. To show the capability of MBRA, we provide several subgoals, specified by latitude, longitude, and azimuth angle values, at intervals of approximately 80 meters, and evaluate LogoNav with MBRA on traversing these subgoals in two scenes. As shown in Fig. 4, our navigation system with our policy enables us to navigate the robot toward a goal 300 meters away without collision, even in human-occupied spaces.
Moreover, we deployed LogoNav on two more robotic embodiments, including VizBot [niwa2022spatio] in an indoor setting, and the Unitree Go1 quadruped robot in an outdoor setting. We conducted 10 trials from up to 100 meters away in different challenging environments with some obstacles for each embodiment and method. We show the quantitative results in Table III and show the robotic behaviors in Fig. 5 and the supplemental videos. We achieve strong goal-reaching behavior with collision avoidance compared to the strongest baseline in Table II, highlighting the policy’s generalization ability. Note that we apply the same policy in Table II and feed the generated actions without any adaptation. Action conversion is internally applied in each robot setup.
Figure 5: Visualization of cross-embodiment analysis.
V-D Training policies on in-the-wild video with MBRA
For Q2, we evaluate the capability of MBRA on different passive data sources, action-free in-the-wild video. We use the MBRA model to generate the action labels for the in-the-wild videos and train the short-horizon visual navigation policy conditioned on goal images, $\{a^{v}_{i}\}_{i=0\ldots N-1}=\pi^{v}(O_{c},O_{g})$. During training, we use the same objective $J_{il}$ to imitate the action labels generated by MBRA. We train three goal image-conditioned policies, one with the GNM dataset alone and another two with GNM + in-the-wild videos with different annotations, visual SLAM [dpvo] and our MBRA model to evaluate how well MBRA enables us to close the embodiment gap between robot and in-the-wild data.
To evaluate the performance in a variety of situations, we collect the topological memories on four indoor trajectories and four outdoor trajectories and deploy the policies with the ERZ. The distance from the initial node to the goal node is between 10.0 m and 31.0 m. As shown in Table IV, the policy trained with the MBRA-annotated in-the-wild video data has an explicit advantage compared to the policy trained only on the GNM dataset. Although the training dataset does not contain the data from the target ERZ, we achieve a high success rate by training with diverse video data. Besides, visual SLAM often fails on videos with fisheye lenses, dense crowds, or few features, leading to worse performance than the others.
V-E Evaluating MBRA for effective crowd-sourced data use
To answer Q3, we compare MBRA and GCP that demonstrated the strongest performance in Table II. We train several relabelers with different data setups for each method and deploy them as the short-horizon navigation policy in the same eight environments and topological memories as in the previous section to more thoroughly explore the capabilities of each of these relabelers.
TABLE VI: Comparison of MBRA and GCP on short-horizon navigation.
Dataset GCP MBRA
@@ V-E Evaluating MBRA for effective crowd-sourced data use
In the final experiment, we aim to assess the generalization capabilities of MBRA model. To this end, we deploy the short-horizon navigation policy on robots in diverse environments across 6 countries: USA, Mexico, China, Mauritius, Costa Rica, and Brazil. In total, we collect 24 topological graphs and evaluate each trajectory. To the best of our knowledge, we are the first to conduct a global evaluation for visual navigation. We evaluate GCP and MBRA model trained with and without the FrodoBots-2k dataset. Findings are summarized in Table V. MBRA model as short-horizon goal image-conditioned navigation policy had better performance for both goal reaching and subgoal coverage than GCP.
VI Conclusion
Limitations: Our MBRA approach for reannotating noisy crowd-sourced data and action-free in-the-wild videos for long-horizon navigation conditioned on a 2D goal pose works well, and we confirmed its applicability to other navigation tasks [hirose2025omnivla], though it still leaves room for improvement. In the model-based approach, we may sometimes generate unreasonable actions because of inaccuracies in the robot model. While we find the model-based approach to generally outperform the imitation-based relabeler (GCP), it does require some strong conditions on the model itself that could prove difficult to translate to more complex tasks like manipulation. A promising direction for future work is developing a more accurate differentiable model that incorporates richer 3D geometry, semantic information, and dynamic object behaviors (e.g., pedestrian motion [hirose2023sacson]), which are essential for robust navigation in cluttered indoor environments. Another important extension is to account not only for goal reaching but also for human preferences—particularly in crowded spaces or settings with strong semantic norms (e.g., avoiding grass when inappropriate). While our model inherits some of these behaviors from human operators in the data, such preferences are not explicitly enforced.